# SportsNGEN: Sustained Generation of Multi-player Sports Gameplay

### v1

@@ Abstract
We present a transformer decoder based model, SportsNGEN, that is trained on sports player and ball tracking sequences that is capable of generating realistic and sustained gameplay. We train and evaluate SportsNGEN on a large database of professional tennis tracking data and demonstrate that by combining the generated simulations with a shot classifier and logic to start and end rallies, the system is capable of simulating an entire tennis match. In addition, a generic version of SportsNGEN can be customized to a specific player by fine-tuning on match data that includes that player. We show that our model is well calibrated and can be used to derive insights for coaches and broadcasters by evaluating counterfactual or what if options. Finally, we show qualitative results indicating the same approach works for football.
*[inlinelist,1]label=(),
Figure 1: Frames from a simulated football match using SportsNGEN. The panels depict a passing sequence involving 3 players. The ball is in the red circle, with an arrow depicting the play that follows. Link to video: https://youtu.be/M0kkKiGVNzk
@@ 1 Introduction
The application of machine learning methods has proven beneficial to many sports applications (Zhao et al., 2023). In particular, sports simulation and analysis can provide valuable insights to sports teams when attempting to understand how small changes to player formation or playing style could impact the next period of play, or their chances of winning (Hauri & Vucetic, 2022; Teranishi et al., 2022; Wang et al., 2023). In addition, realistic gameplay simulation is critical in computer gaming scenarios (Kurach et al., 2020).
@@
Tremendous progress has been made in the area of sports trajectory prediction (Yue et al., 2014; Zheng et al., 2016; Le et al., 2017b; Zhan et al., 2019; Li et al., 2021; Tang et al., 2021; Wu et al., 2021; Alcorn & Nguyen, 2021; Omidshafiei et al., 2022), however it is difficult to precisely mimic training data over long periods of time. Figure A.1 shows how the prediction error of the player and ball positions increases with time when simulated tennis data from our system is compared to the training data. Sports are inherently unpredictable over longer time scales and so deterministic prediction is not possible or useful in many scenarios. Instead, it is important to capture different ways the match will evolve in a statistically correct way.
Significant advancements have been made in sports simulation by leveraging reinforcement learning (RL) techniques (Kurach et al., 2020; Liu et al., 2021; Braga & Barros, 2022; Yu et al., 2023). Recently, the transformer architecture (Vaswani et al., 2017) has been applied to multi-agent spatiotemporal systems problems in order to generate realistic sports simulations and understand player behavioural patterns (Alcorn & Nguyen, 2021). Instead of generating words as in natural language processing, player and ball movements over time can be generated by training a transformer model to predict the next position from a sequence of player tracking data.
@@
We propose that generated sports simulations should be: 1 highly realistic both visually and statistically similar to real gameplay data; 2 sustainedfor the duration between natural breaks in the gameplay; 3 customizablevia fine-tuning or other method to emulate the style of play of a particular player and/or team; and 4 measureablein that metrics are available to evaluate the quality of the simulations (as opposed relying on a human expert) such that the simulations can be improved by optimizing the metrics.
@@
However, to the best of our knowledge, no previous work has been successful in generating realistic, sustained, and customizable simulations learned from player and ball tracking data for more than short periods of time. In this work we present Sports Neural Generator or SportsNGEN that realizes the goals of realistic, sustained, customizable and measurable sports gameplay. Figure 1 and Figure 2 depict football [^1] We use the term football to refer European football or soccer. and tennis sequences, respectively, generated by our approach along with links to simulation videos.
@@
Our contributions: 1 A transformer decoder based model, SportsNGEN, trained on player and ball tracking data as well as match metadata that is capable of generating gameplay simulations that are statistically similar to the training data and are sustained for the duration of normal breaks in play. 2 We demonstrate through ablations that the following enhancements significantly improve generated simulations: a) extending the player and ball representations to include relative velocity, distance to the ball, and time into the game or sequence; and b) adding small perturbations to the ball positions during training to allow the model to correct for errors. 3 Training and evaluating SportsNGEN on a large database of professional tennis tracking data that is capable of simulating an entire tennis match by combining the generated simulations with a shot classifier and logic to start and end rallies. 4 We introduce metrics to statistically evaluate the quality of generated tennis gameplay. 5 We demonstrate that a generic version of our model can be customized to a specific tennis player by fine-tuning on match data that includes that player. 6 Finally, we show that our model can be used to inform coaching decisions by evaluating counterfactual or what if options.
@@ 2 Related Work
In this section, we discuss related work in the categories of sports analytics, and game simulation. See Section A.11 for related work pertaining to trajectory prediction.
@@
Group Activity Recognition and Sports Analytics Miller et al. (2014) develop an approach to represent and analyze the underlying spatial structure that governs shot selection among professional basketball players. Le et al. (2017a) employ an imitation learning approach to analyze football defensive strategies. Hauri & Vucetic (2022) propose a transformer-based architecture with a Long Short-Term Memory (LSTM) embedding to recognize basketball group activities from player and ball tracking data. Teranishi et al. (2022) evaluate football players who create off-ball scoring opportunities by comparing actual movements with the reference movements generated via trajectory prediction. Chen et al. (2023) use a probabilistic diffusion approach to model basketball player behavior. The model only considers player movement and no other metadata. Wang et al. (2023) present a football tactics assistant that focuses on analyzing corner kicks which allows coaches to explore player setup options and use those with the highest likelihood of success.
@@
Game Simulation Kurach et al. (2020) introduce a game engine that simulates football gameplay with an environment for evaluating RL algorithms. Liu et al. (2021) demonstrate an RL approach, where the agents progressively learn to play football initially from random behavior, to simple ball chasing, to showing evidence of cooperation. Braga & Barros (2022) introduce a simulator for robot football optimized for performing RL experiments. Yu et al. (2023) introduce a RL environment where agents are trained to play basketball.
@@
Our approach is not RL based. It instead learns in a discriminative fashion from sequences of gameplay tracking data, which obviates the need to use physics based models of gameplay or learning gameplay from scratch with RL. This also enables us to build predictive models for specific players which can be important for analysis and gaming scenarios. Overall, our work is distinct from the above works in that our goal is to generate sports gameplay for the entire duration of a match that is statistically similar to the training data.
@@ 3.1 Input Data
3.2 Transfer Decoder Model
We use a transformer decoder model $f$ that is an extended implementation of baller2vec (Alcorn & Nguyen, 2021) to predict future player and ball states given the current and recent history of states. The model $f$ is run in an auto-regressive mode with a rolling window of length $T$, using a specified period of previous predictions to predict the ball and player state at the next step. We use the same attention method as baller2vec, permitting each object token to attend to every object token up to and including its own time step. We adjust the attention mask so that each object token can attend to the context tokens, influencing the predictions for player and ball movement. We treat the update step as a classification as opposed to a regression or diffusion problem, by splitting the area of possible next locations for the ball and players into a 3D and 2D grid, respectively, of discrete bins that indicate the relative offset $\rho$ from the current position $p$ as this is easier to learn and can bound motion to physically possible values. A depiction of a grid for a football player and the ball is shown in Figure 5.
@@ 3.2 Transfer Decoder Model
We use nucleus sampling (Holtzman et al., 2020) to sample the location in the output grid based on the output probabilities of $f$. When the grid location has been selected, we turn the discrete value into a continuous value by sampling from a uniform distribution across the bin. If the initial conditions for the player or ball have zero velocity, this helps to force the simulation into motion by avoiding continuous velocity predictions of zero. To enable the model to learn the behavior of individual players, the bin size must be fine grained enough for predictions to capture distinguishing features. In many sports, important statistics include how fast a player can run, or how far they can hit, throw or kick the ball. Formally, the probability distribution of predicting a particular bin location $k$ for an object $n$ at step $\tau+1$ is
@@
Based on the sampled value of $\rho$ and the mapping between bins and physical distance, the updated values of position $p_{\tau+1,n}$, velocity $v_{\tau+1,n}$, and distance to the ball $d_{\tau+1,n}$ can be computed. Since we use the baller2vec attention mask, the positions of the ball and each player can be updated simultaneously at each time step. We detect the end of a simulation or break in a play with logic specific to each sport. For example, we can end simulations if a ball goes out of bounds or in some sports if the ball makes contact with the ground, or if the time in the period of play runs out. When generating simulations, we set a maximum input sequence length of $T$ tokens. For a player and ball state update at step $\tau+1$, we input from $\tau-T$ to $\tau$ steps of initial token data into the model $f$. If $T$ time steps of data are not yet simulated, the missing tokens are padded with zeros and masked. Specifically, simulations are rolled autoregressively out at the $i$th step as
@@
3.3 Event Classification and Transfer Learning
@@
As an extension to training a model capable of capturing the behavior of all players, we also train a generic model $f_{gen}$ which learns a single feature vector $I_{gen}$, called the generic player vector where $I_{n}=I_{gen},n\in N$. We then fine-tune $f_{gen}$ with matches containing a specific player or team, and transfer learn a new set of $I_{n}\in N$ for that player or team that can represent their behavior against a generic opponent.
@@ 4 Tennis Implementation Details
We use a proprietary dataset of tennis tracking data for approximately 15,000 tennis matches containing 7.6 million rally sequences. The data contain COM locations for each player and the ball sampled at 25 Hz, with the center of the court at $(x,y,z)=(0,0,0)$, whose components refer to the length, width, and vertical directions, respectively. The data also contain metadata about each match and rally, containing: the players in the rally, the tournament and court, the rally winner, whether the rally was a first or second serve, and what shots were played. The tracking data set is cut up into individual sequences that start at the toss before a serve and end shortly after the rally is finished. We set a maximum sequence length of input data to be 6 seconds. We found that increasing the sequence length to be more than 6 seconds became computational impractical and did not improve the model accuracy. We also double the size of the data set by flipping the data along the $x$ and $y$ axes.
@@
The playing surface is important contextual information when predicting rallies in tennis. The expectation is that hard and grass courts have the fastest bounces, and clay courts absorb more momentum from the impact resulting in slightly slower bounces. We learn context vectors for each surface and tournament in the dataset, and also encourage the model to learn the difference between first and second serve types by including context vectors for both.
@@
To understand who won the rally, and for analysis of the point, we train the event classifier to classify the type of shot being played at every step within the simulation. This includes the type of stroke (groundstroke, serve, volley, etc.), the direction of the shot (cross court, down the line, etc.), whether the shot is a winner, error or a continuation of the point, and if an error is forced or unforced. The event classifier $g$ receives as input a simplified version of the input token, without any identity $I$ or context $C$ components. In the training data, shot type labels are consistent across time steps between shots. The model is expected to predict the same, only varying its prediction when the ball contacts a racket. When a rally is finished, we convert the tracking data from the rally into the shot type classifier input, run the model once to identify where the changes in shot type are, and take the model shot type between changes as the final label for each shot. The winner or error classification for the final shot of the rally tells us who won the point, and the shot type labels help us break down the shots for statistical analysis. To combine rallies together to simulate an entire match, all that is left to do is implement logic to increment the score, calculate who is serving, from which end and which side. These can be used to obtain the initial conditions for the next point. Figure A.9 shows the validation loss for the shot classifier for each event. The low held out loss values indicate that it functions as intended.
Section A.2 details the network architecture for the tennis implementation of the transformer decoder $f$.
@@ 5 Experiments
For the tennis experiments, we selected 3 male professional players with varying styles to evaluate SportsNGEN and simulated 6 matches between each combination of two players, each match was the best of 3 sets. We repeat this experiment across 3 different tournaments, one for each surface type: hard, clay and grass. For comparison, we then collect data from the training data set, where these players have played each other on these surfaces. Using both real and simulated data, we compute relevant statistics and define an evaluation metric for each statistic as the difference between the two.
For physical metrics, we compare the median, inter-quartile range (IQR), and Wasserstein distance between the distributions of real and simulated data for the following quantities collected across all matches: 1 Toss contact height: Height of the ball at the contact point with the racket during serving. 2 First and second serve speeds: Maximum recorded speed during the serve. 3 Return speeds: Maximum speed of a return of serve. 4 Groundstroke speeds: Maximum speed of all groundstrokes.
@@
Varying the top-$p$ sampling parameter Figure 6 shows typical output probability distributions, projected into the $xy$ plane, for an update step of the ball in mid-flight, and at the moment the ball is about to be hit. The peaks in intensity for the mid-flight predictions (a) are distributed over very few bins since the model has learned the physical constraints of the system (e.g. drag, gravity), and can therefore be very confident in how to update the ball state. The remaining panels (b)-(d) depict the probability distributions for the ball at the time of impact (ToI) – the point at which a player hits the ball for various values of top-$p$. The distributions in these cases contain multiple separated peaks in intensity in the $xy$ plane. This corresponds to the choice to play the shot either down the line or across the court which enables us to perform counterfactual analysis (see Section 5). As top-$p$ decreases, the probability of a cross court shot decreases. In general we will see that a low top-$p$ value will result in less variety in playing style.
@@
Calibration A key intended application of SportsNGEN is generating insights for coaching and sports broadcasts. For these applications the model should accurately forecast the probability that each player wins a rally as it develops. We can test SportsNGEN’s ability to do this in the following way. We sample random rallies from the training data, and roll out the model from a given random time step 100 times, to generate a win percentage for both players. Repeating this for a large number of starting points, we form a histogram of predictions by stratifying the predictions into bins (Figure A.4). For each prediction, we also have the ground truth of who won the rally in the training data. So, taking the 90% bin for example, if the model is well-calibrated, the corresponding ground truth rallies should be won by the player in 90% of cases. Figure 9 shows that the win percentages generated by the SportsNGEN are well-calibrated, with deviations where data are sparse.
[…]
Context Token Study We add context tokens to encode the tournament, court surface type, and whether the serve is the player’s first or second. Typically the second serve is expected to be slower since players will prioritize accuracy over speed to avoid losing a point through double fault. Figure 10 shows that the addition of a serve context token $C_{serve}$ as well as the player ID component $I$ in $O$ reduce the difference between real and simulated serve speeds and produce narrower distributions between first and second serve speeds. The results for additional players are shown in Figure A.7.
Figure 10: The distribution of first and second serve speeds. a) Using a serve context token $C_{serve}$ and the ID component $I$ in $O$; b) Using $C_{serve}$ but not $I$; c) Using $I$ but not $C_{serve}$. The vertical lines show the real and simulated median serve speeds.
@@
Transfer Learning LABEL:png shows various metrics as a function of the number of training sequences that are required to fine-tune $f_{gen}$ such that the generic player ID vector $I$ is adapted to a new player. In the simulations, $f_{gen}$ is the opponent for the fine-tuned model. The groundstroke and return metrics improve as the number of training samples increases whereas the serve metrics fluctuate with the first serve speed getting worse. This can be explained by the low variability of the serve distribution being easier to learn when compared to highly variable groundstroke patterns.
Training Time Analysis Section A.3 shows the various metrics as a function of training iteration. All metrics improve as the number of training iterations are increased although there are diminishing returns after 80000.

### v3

# SportsNGEN: Sustained Generation of Realistic Multi-player Sports Gameplay

@@ Abstract
We present a transformer decoder based sports simulation engine, SportsNGEN, trained on sports player and ball tracking sequences, that is capable of generating sustained gameplay and accurately mimicking the decision making of real players. By training on a large database of professional tennis tracking data, we demonstrate that simulations produced by SportsNGEN can be used to predict the outcomes of rallies, determine the best shot choices at any point, and evaluate counterfactual or what if scenarios to inform coaching decisions and elevate broadcast coverage. By combining the generated simulations with a shot classifier and logic to start and end rallies, the system is capable of simulating an entire tennis match. We evaluate SportsNGEN by comparing statistics of the simulations with those of real matches between the same players. We show that the model output sampling parameters are crucial to simulation realism and that SportsNGEN is probabilistically well-calibrated to real data. In addition, a generic version of SportsNGEN can be customized to a specific player by fine-tuning on the subset of match data that includes that player. Finally, we show qualitative results indicating the same approach works for football.
Figure 1: Frames from a football match simulated using SportsNGEN. The panels depict a passing sequence involving 3 players. The ball is in the red circle, with an arrow depicting the play that follows. Link to video: https://youtu.be/M0kkKiGVNzk
@@ 1 Introduction
The application of machine learning methods has proven beneficial to many sports applications (Zhao et al.,, 2023). In particular, sports simulation and analysis can provide valuable insights to sports teams when attempting to understand how small changes to player formation or playing style could impact the next period of play, or their chances of winning (Hauri and Vucetic,, 2022; Teranishi et al.,, 2022; Wang et al.,, 2023). In addition, realistic gameplay simulation is critical in computer gaming scenarios (Kurach et al.,, 2020).
@@
Tremendous progress has been made in the area of sports trajectory prediction (Yue et al.,, 2014; Zheng et al.,, 2016; Le et al., 2017b, ; Zhan et al.,, 2019; Li et al.,, 2021; Tang et al.,, 2021; Wu et al.,, 2021; Alcorn and Nguyen,, 2021; Omidshafiei et al.,, 2022), however it is difficult to precisely mimic training data over long periods of time. Figure 18 shows how the prediction error of the player and ball positions increases with time when simulated tennis data from our system is compared to the training data. Sports are inherently unpredictable over time scales longer than several seconds and as a result, deterministic prediction is not possible or useful in many scenarios. Instead, it is important to capture the different ways a match will evolve in a statistically accurate manner by modelling the complete distribution of player decision making.
@@
We propose that generated sports simulations should be: 1 highly realistic and capture the complete distribution of real player behaviour; 2 sustainedfor the duration between natural breaks in the gameplay; 3 customizablevia fine-tuning or other method to emulate the style of play of a particular player and/or team; and 4 measurablein that metrics are available to evaluate the quality of the simulations (as opposed relying on a human expert) such that the simulations can be improved by optimizing the metrics.
Recently, the transformer architecture (Vaswani et al.,, 2017) has been applied to multi-agent spatiotemporal systems problems to generate realistic sports simulations and understand player behavioural patterns (Alcorn and Nguyen,, 2021). Instead of generating words as in natural language processing, temporal player and ball movements can be generated by training a transformer model to predict the next position from a sequence of tracking data.
@@
However, to the best of our knowledge, no previous work has been successful in generating realistic, sustained, and customizable simulations, learned from player and ball tracking data, for more than a few seconds. In this work we present Sports Neural Generator or SportsNGEN that realizes the goals of realistic, sustained, customizable and measurable sports gameplay. Figure 1 and Figure 2 depict football [^1] We use the term football to refer European football or soccer. and tennis sequences, respectively, generated by our approach along with links to simulation videos.
@@
Our contributions: 1 A transformer decoder based simulation engine, SportsNGEN, trained on player and ball tracking data as well as match metadata, capable of simulating the movement of all players and the ball simultaneously in a sports game scenario. The simulations are sustained between breaks in play. 2 Training and evaluating SportsNGEN on a large database of professional tennis tracking data. SportsNGEN is capable of simulating an entire tennis match by combining the generated simulations with a shot classifier and logic to start and end rallies. 3 We show that our model can be used to inform tennis coaching decisions and best shot options by evaluating counterfactual or what if options. 4 We demonstrate through ablations that the following enhancements significantly improve convergence and generated simulations: a) extending the player and ball representations to include relative velocity, distance to the ball, and time into the game or sequence; b) adding small perturbations to the ball positions during training to allow the model to correct for errors; and c) Adding context tokens to allow the model to adapt to different playing surfaces. 5 We devise a novel optimization method by defining metrics to statistically evaluate the quality of generated tennis data. By altering simulation hyperparameters, we show that the simulations can be optimized to be statistically similar to the behaviour of real players. 6 We demonstrate that a generic version of our model can be customized to a specific tennis player by fine-tuning on match data that includes that player.
@@ 2 Related Work
In this section, we discuss related work in the categories of sports analytics, and game simulation, and trajectory prediction.
Group Activity Recognition and Sports Analytics
@@
Miller et al., (2014) develop an approach to represent and analyze the underlying spatial structure that governs shot selection among professional basketball players. Le et al., 2017a employ an imitation learning approach to analyze football defensive strategies. Hauri and Vucetic, (2022) propose a transformer-based architecture with a Long Short-Term Memory (LSTM) embedding to recognize basketball group activities from player and ball tracking data. Teranishi et al., (2022) evaluate football players who create off-ball scoring opportunities by comparing actual movements with the reference movements generated via trajectory prediction. Chen et al., (2023) use a probabilistic diffusion approach to model basketball player behavior. The model only considers player movement and no other metadata. Wang et al., (2023) present a football tactics assistant that focuses on analyzing corner kicks which allows coaches to explore player setup options and use those with the highest likelihood of success.
Game Simulation
@@
Kurach et al., (2020) introduce a game engine that simulates football gameplay with an environment for evaluating RL algorithms. Liu et al., (2021) demonstrate an RL approach, where the agents progressively learn to play football initially from random behavior, to simple ball chasing, to showing evidence of cooperation. Braga and Barros, (2022) introduce a simulator for robot football optimized for performing RL experiments. Yu et al., (2023) introduce a RL environment where agents are trained to play basketball.
Finally, Yuan et al., (2023) describe a method to learn simulated tennis skills from broadcast videos. However, this approach only models one shot cycle at a time, using statistical analysis to predict the desired shot location for specific players. There is no coupling between the current shot and previous shots, so strategic play is limited. Also, the players are only allowed to move on the baseline, so no volleys or inner court play is permitted, restricting realism.
@@
Our approach is not RL based. It instead learns in a discriminative fashion from sequences of gameplay tracking data, which obviates the need to use physics based models of gameplay or learning gameplay from scratch with RL. This also enables us to build predictive models for specific players which can be important for analysis and gaming scenarios. Overall, our work is distinct from the above works in that our goal is to generate sports gameplay that captures the complete distribution of player behaviour, where the aggregation of non-deterministic simulations is statistically similar to real data. This acts as a powerful tool for strategic analysis, evaluating how player decision making affects the outcome of a period of play.
Sports Trajectory Prediction
There is a rich literature on trajectory prediction in general, and sports trajectory prediction in particular. Yue et al., (2014) learn predictive models for basketball play prediction given the current game state. Zheng et al., (2016) model spatiotemporal trajectories over long time horizons using expert demonstrations capable of generating realistic, but short rollouts. Le et al., 2017b present an LSTM based imitation learning approach for learning multiple policies for team defense in professional football. However, no policy is learned for the position of the ball. Zhan et al., (2019) describe a hierarchical framework for sequential generative modeling that can generate high quality trajectories and encode coordination between agents. However, their framework cannot generate entire games. Li et al., (2021) describes an approach for multi-agent trajectory prediction using a graph neural network. When evaluated on basketball data, only short trajectories were considered. Tang et al., (2021) propose the concept of collaborative uncertainty, to model the uncertainty in interaction in multi-agent trajectory forecasting. Wu et al., (2021) propose a generative adversarial network (GAN) to generate short basketball player and ball trajectories. Alcorn and Nguyen, (2021) introduce baller2vec, a multi-entity transformer that can model coordinated agents. It employs a special self-attention mask to learn the distributions of statistically dependent agent trajectories and is shown to generate realistic trajectories for basketball players or the ball itself for short durations. Our work builds upon baller2vec to enable sustained gameplay simulations by simultaneously simulating both the player and the ball. Omidshafiei et al., (2022) study the problem of multiagent time-series imputation in the context of football in order to predict the behaviors of off-screen players.
@@ 3.1 Input Data
3.2 Transformer Decoder Model
We use a transformer decoder model $f$ that is an extended implementation of baller2vec (Alcorn and Nguyen,, 2021) to predict future player and ball states given the current and recent history of states. We make a significant update to the baller2vec experiments by modelling both the ball and the players simultaneously.
The model $f$ is run in an auto-regressive mode with a rolling window of length $T$, using a specified period of previous predictions to predict the ball and player state at the next step. We use the same attention method as baller2vec, permitting each object token to attend to every object token up to and including its own time step. We adjust the attention mask so that each object token can attend to the context tokens, influencing the predictions for player and ball movement.
We treat the update step as a classification as opposed to a regression or diffusion problem, by splitting the area of possible next locations for the ball and players into a 3D and 2D grid, respectively, of discrete bins that indicate the relative offset $\rho$ from the current position $p$ as this is easier to learn and can bound motion to physically possible values. A depiction of a grid for a football player and the ball is shown in Figure 5.
@@ 3.2 Transformer Decoder Model
We use nucleus sampling (Holtzman et al.,, 2020) to sample the location in the output grid based on the output probabilities of $f$. When the grid location has been selected, we turn the discrete value into a continuous value by sampling from a uniform distribution across the bin. If the initial conditions for the player or ball have zero velocity, this helps to force the simulation into motion by avoiding continuous velocity predictions of zero.
To enable the model to learn the behavior of individual players, the bin size must be fine grained enough for predictions to capture distinguishing features. In many sports, important statistics include how fast a player can run, or how far they can hit, throw or kick the ball. Formally, the probability distribution of predicting a particular bin location $k$ for an object $n$ at step $\tau+1$ is
@@
Based on the sampled value of $\rho$ and the mapping between bins and physical distance, the updated values of position $p_{\tau+1,n}$, velocity $v_{\tau+1,n}$, and distance to the ball $d_{\tau+1,n}$ can be computed. Since we use the baller2vec attention mask, the positions of the ball and each player can be updated simultaneously at each time step.
We detect the end of a simulation or break in a play with logic specific to each sport. For example, we can end simulations if a ball goes out of bounds or in some sports if the ball makes contact with the ground, or if the time in the period of play runs out. When generating simulations, we set a maximum input sequence length of $T$ tokens. For a player and ball state update at step $\tau+1$, we input from $\tau-T$ to $\tau$ steps of initial token data into the model $f$. If $T$ time steps of data are not yet simulated, the missing tokens are padded with zeros and masked. Specifically, simulations are rolled autoregressively out at the $i$th step as
@@
3.3 Event Classification
@@ 4 Tennis Implementation Details
We use a proprietary dataset of tennis tracking data. The data contains COM locations for each player and the ball sampled at 25 Hz, with the center of the court at $(x,y,z)=(0,0,0)$, whose components refer to the length, width, and vertical directions, respectively.
The data also contains metadata about each match and rally, including: the players in the rally, the tournament and court, the rally winner, whether the rally was a first or second serve, and what shots were played. The tracking data set is cut up into individual sequences that start at the toss before a serve and end shortly after the rally is finished. We set a maximum sequence length of input data to be 6 seconds. We found that increasing the sequence length to be more than 6 seconds became computational impractical and did not improve the model accuracy. This suggests that professional tennis players’ decision making is not strongly affected by information further than 6 seconds into the past. We also double the size of the data set by flipping the data along the $x$ and $y$ axes simultaneously.
@@
The playing surface is important contextual information when predicting rallies in tennis. The expectation is that hard and grass courts have the fastest bounces, and clay courts absorb more momentum from the impact resulting in slightly slower bounces and longer rallies. We learn context vectors for each surface and tournament in the dataset, and also encourage the model to learn the difference between first and second serve types by including context vectors for both.
@@
To understand who won the rally, and for analysis of the point, we train the event classifier to classify the type of shot being played at every step within the simulation. This includes the type of stroke (groundstroke, serve, volley, etc.), the direction of the shot (cross court, down the line, etc.), whether the shot is a winner, error, or a continuation of the point, and if an error is forced or unforced.
The event classifier $g$ receives as input a simplified version of the input token, without any identity $I$ or context $C$ components. In the training data, shot type labels are consistent across time steps between shots. The model is expected to predict the same, only varying its prediction when the ball contacts a racket. When a rally is finished, we convert the tracking data from the rally into the shot type classifier input, run the model once to identify where the changes in shot type are, and take the model shot type between changes as the final label for each shot. The winner or error classification for the final shot of the rally tells us who won the point, and the shot type labels help us break down the shots for statistical analysis. To combine rallies together to simulate an entire match, all that is left to do is implement logic to increment the score, calculate who is serving, from which end and which side. These can be used to obtain the initial conditions for the next point.
Tennis Network Architecture
The input tokens $O_{\tau,n}$ are embedded with a 3 layer MLP with input size 30, hidden sizes 256 and 512, and output size 2048. The transformer decoder, $f$, has 4 layers, 2048 embedding dimension, 8 heads, 4 expansion factor, and 0.2 dropout. The shared player output network is a single linear layer with input size 2048 output size equal to 61 $\times$ 61 bins. The ball output network is a single linear layer with input size 2048 and output size equal to 61 $\times$ 61 $\times$ 61 bins.
@@ 5 Experiments
For the tennis experiments, we selected 3 male professional players with varying styles to evaluate SportsNGEN and simulated 6 matches between each combination of two players. Each match was the best of 3 sets. We repeat this experiment across 3 different tournaments, one for each surface type: hard, clay and grass. For comparison, we then collect data where these players have played each other on these surfaces. Using both real and simulated data, we compute relevant statistics and define an evaluation metric for each statistic as the difference between the two.
For physical metrics, we compare the median, inter-quartile range (IQR), and Wasserstein distance between the distributions of real and simulated data for the following quantities collected across all matches:
  • Toss contact height: Height of the ball at the contact point with the racket during serving.
  • First and second serve speeds: Maximum recorded speed during the serve.
  • Return speeds: Maximum speed of a return of serve.
  • Groundstroke speeds: Maximum speed of all groundstrokes.
We also compute additional relevant statistics based on aggregated data. For these quantities, a scalar value is aggregated over many rallies for each player. The absolute difference between the real and simulated aggregated scalars is compared.
[…]
@@
Varying the top-$p$ sampling parameter
Figure 6 shows typical output probability distributions, projected into the $xy$ plane, for an update step of the ball in mid-flight, and at the moment the ball is about to be hit. The peaks in intensity for the mid-flight predictions (a) are distributed over very few bins since the model has learned the physical constraints of the system (e.g. drag, gravity), and can therefore be very confident in how to update the ball state. The remaining panels (b)-(d) depict the probability distributions for the ball at the time of impact (ToI) – the point at which a player hits the ball for various values of top-$p$. The distributions in these cases contain multiple separated peaks in intensity in the $xy$ plane. This corresponds to different decisions a player may make when choosing to play the shot either down the line or across the court. By sampling from these two modes, we are able to perform counterfactual analysis (see Section 5). As top-$p$ decreases, the probability of sampling a cross court shot decreases, demonstrating the need to optimize top-$p$ in order to accurately capture the player behaviour in the simulations. In general, we will see that a low top-$p$ value will result in less variety in playing style, but a high top-$p$ value will result in many outliers.
@@
Figure 8: A comparison of convergence for SportsNGEN against a Baseline model without $v$, $d$, $e$ and context tokens $C$, for a) An average of the 4 physical metrics, and b) Non-realistic rallies as a function of training iterations.
Figure 7b) and (c) show how the various metrics vary with top-$p$. In (b), the number of non-realistic rallies (rallies that must be discarded based on logical checks) increases with a value of top-$p$ both that is too high, and too low. Increasing top-$p$ increases the probability that the ball trajectory may be updated in a way that defies physical constraints and would be forced to be removed. With too low top-$p$ there may be too few options for the ball and player to update in a way that leads to a realistic rally. In (c), we see that with the exception of double fault percentage, the metrics reach optimal values when top-$p$ is in the range of 0.8 to 0.9.
Object Token Component Ablation Study
@@
Varying the size $\iota$ of the player encoding vector $I$ in Figure 9, we find that the accuracy increases until $\iota=$20 where there are diminishing returns for further increases. This is also supported by Figure 10 d)-f), where the data with no player ID $I$ has a much broader distribution of serve speeds, and a nearly identical median serve speed for all three players.
Figure 9: Varying the player ID $I$ size $\iota$ to show how various metrics can be improved with a larger $\iota$. As a control, we train a generic model without player ID $I$.
Context Token Study
We add context tokens to encode the tournament, court surface type, and whether the serve is the player’s first or second. Typically the second serve is expected to be slower since players will prioritize accuracy over speed to avoid losing a point through double fault. Figure 10 shows that the addition of a serve context token $C_{serve}$ as well as the player ID component $I$ in $O$ reduce the difference between real and simulated serve speeds and produce narrower distributions between first and second serve speeds.
Figure 10: The distribution of first and second serve speeds for all three players for the following models: (top) SportsNGEN, (middle) a model with no player ID vector $I$, (bottom) a model with no serve context token $C_{serve}$.
@@
As an extension to training a model capable of capturing the behavior of all players, we also train a generic model $f_{gen}$ which learns a single feature vector $I_{gen}$, called the generic player vector where $I_{n}=I_{gen},n\in N$. We then fine-tune $f_{gen}$ with matches containing a specific player, and transfer learn a new set of $I_{n}\in N$ for the player that can represent their behavior against a generic opponent. This could be used for quickly customizing a pretrained model to a new player on the circuit.
Figure 13 shows various metrics as a function of the number of training sequences that are required to fine-tune $f_{gen}$ such that the generic player ID vector $I$ is adapted to a new player. In the simulations, $f_{gen}$ is the opponent for the fine-tuned model. The groundstroke and return metrics improve as the number of training samples increases whereas the serve metrics fluctuate with the first serve speed getting worse. This can be explained by the low variability of the serve distribution being easier to learn when compared to highly variable groundstroke patterns.